OpenStudy (anonymous):

A deck of 320 punched cards can be read by one reader in 20 minutes. A second reader can read the same deck in 12 minutes. How many minutes would it take both readers working together to read a deck? How long would it take to process the deck of cards in example one if the first reader stopped working after 5 minutes?

OpenStudy (anonymous):

Denote \(R_1, R_2\) to be the average reading speed of each reader, then, we have: \[ R_1=\frac{320}{20}=16\\ R_2=\frac{320}{12}=\;\frac{80}{3} \]So, if both work in perfect parallel motion, we have that their combined reading speed is: \[ 16+\frac{80}{3}=\frac{128}{3} \]So, it would take a total of: \[ \frac{320}{\frac{128}{3}}=320\cdot\frac{3}{128}=7.5\text{ minutes} \]To read the whole deck.

OpenStudy (anonymous):

If one of them stopped working after 5 minutes, then, we have that they would have read a combined total of: \[ 16\cdot5+\frac{80}{3}\cdot5=80+\frac{400}{3}=213+\frac{1}{3}\text{ cards} \]So, we get a total of \[ 320-(213+\frac{1}{3})=146+\frac{2}{3} \]Remaining cards. We only have the second machine working, so we get: \[ \frac{146+\frac{2}{3}}{\frac{128}{3}}=(146+\frac{2}{3})\cdot\frac{3}{128}=3.4375\text{ minutes} \]To finish the deck, giving us a combined total of: \[ 5+3.4375=8.4375\text{ minutes} \] Wow... I sure hope I didn't screw up any of the calculations.

OpenStudy (anonymous):

Wait, wait, wait, wait, yes I did... For that last part, it would take: \[ \frac{146+\frac{2}{3}}{\frac{80}{3}}=5.5\text{ minutes} \]More, so, we get that it would take: \[ 5.5+5=10.5\text{ minutes} \]